What is the frequency difference of 100 GHz equivalent to at the 850nm wavelength of Gallium arsenide lasers?
1
Expert's answer
2012-11-12T11:45:51-0500
850 nm transformed into Hz is
c/nu = 3*10^8/850*10^(-9) = 3.52*10^(14) Hz

100 GHz = 10^(11) Hz is much less then 850 nm in Hertz.